OpenStudy (2yoututu11):

The number of traffic violations in a city increased from 12 per month to 16 per month over the course of two years. What is the percent of increase in the number of traffic violations per month for that two-year period? Round to the nearest percent, if necessary.

OpenStudy (anonymous):

question is worded a bit dodgy 12 to 16 is an increase of 33.3%

OpenStudy (anonymous):

Increase is from 12 to 16 so an increase of 4. As a percentege of 12 the calculation is:- 4/12 x 100 = 400/12 = 33.33 recurring so to nearest percent would be 33%

OpenStudy (2yoututu11):